### v1

# Uncovering Treatment Effect Heterogeneity in Pragmatic Gerontology Trials

Abstract
Detecting heterogeneity in treatment response enriches the interpretation of gerontologic trials. In aging research, estimating the effect of the intervention on clinically meaningful outcomes faces analytical challenges when it is truncated by death. For example, in the Whole Systems Demonstrator trial, a large cluster-randomized study evaluating telecare among older adults, the overall effect of the intervention on quality of life was found to be null. However, this marginal intervention estimate obscures potential heterogeneity of individuals responding to the intervention, particularly among those who survive to the end of follow-up. To explore this heterogeneity, we adopt a causal framework grounded in principal stratification, targeting the Survivor Average Causal Effect (SACE)—the treatment effect among “always-survivors,” or those who would survive regardless of treatment assignment. We extend this framework using Bayesian Additive Regression Trees (BART), a nonparametric machine learning method, to flexibly model both latent principal strata and stratum-specific potential outcomes. This enables the estimation of the Conditional SACE (CSACE), allowing us to uncover variation in treatment effects across subgroups defined by baseline characteristics. Our analysis reveals that despite the null average effect, some subgroups experience distinct quality of life benefits (or lack thereof) from telecare, highlighting opportunities for more personalized intervention strategies. This study demonstrates how embedding machine learning methods, such as BART, within a principled causal inference framework can offer deeper insights into trial data with complex features including truncation by death and clustering—key considerations in analyzing pragmatic gerontology trials.
Keywords: Whole Systems Demonstrator Telecare Questionnaire Study; gerontology trials, cluster randomized trial; pragmatic trials, principal stratification; machine learning; Bayesian inference; survivor average causal effect
1 Introduction
Pragmatic trials are a cornerstone of gerontologic research, providing rigorous yet policy-relevant evidence on interventions that aim to support independence, quality of life, and health in late life.roland1998understanding; resnick2022pragmatic Their strengths lie in large, heterogeneous populations and real-world delivery, often using cluster-randomized designs. eldridge2012practical However, these very characteristics can complicate the inference on the intervention performance: treatment effects averaged across diverse participants may appear modest or null, even when meaningful benefits (or harms) exist for specific subgroups. Conventional subgroup analyses, typically prespecified along a few baseline dimensions, are rarely sufficient to detect more complex forms of heterogeneity. As a result, valuable information may remain hidden, and seemingly “negative” trials risk being dismissed despite containing signals that could guide intervention targeting and future research. As a result, machine learning and other data-driven approaches are increasingly valuable for estimating individualized treatment effects and detecting complex, multidimensional patterns of heterogeneity that conventional subgroup analyses may overlook kent2018personalized; halvorsrud2007conceptualization. As one prominent example, parallel advances in machine learning for causal inference have shown that Bayesian Additive Regression Trees (BART) can flexibly learn complex regression surfaces with built-in regularization chipman2010bart, and have been recently applied to treatment effect heterogeneity detection in clinical trials. chen2024bayesian; hu2021estimating; ghazi2025treatment.
Yet, in gerontologic trials, these opportunities come with unique complications. quality of life (QoL) outcomes that are among the most clinically meaningful endpoints in late-life intervention studies are a prime example. halvorsrud2007conceptualization; netuveli2008quality Measures such as EQ-5D or SF-12 are widely used to capture well-being, functional status, and independence—dimensions of aging that survival alone cannot reflect. group1990euroqol; ware199612 However, QoL outcomes face a special challenge: they are undefined for participants who die before follow-up, or who deteriorate so severely that reporting is impossible. This phenomenon, known as truncation by death, is especially common in gerontology trials, where mortality is non-negligible during extended follow-up. rubin2006causal Because participants who die may differ systematically from those who survive, and because interventions themselves may influence participant survival, analyses restricted to observed survivors risk introducing survivor bias and yield estimates that are difficult to interpret causally. One principled solution comes from principal stratification, which defines causal effects within strata of participants classified by their potential survival under both treatment and control.frangakis2002principal; zhang2003estimation In this framework, the survivor average causal effect (SACE) focuses on the “always-survivors”—those who would survive regardless of assignment—for whom both counterfactual QoL outcomes are well defined. Extending this idea to covariate conditional effects (CSACE) enables the investigation of how survivor responses vary across baseline characteristics, uncovering heterogeneity in a way that standard subgroup analyses cannot.chen2024bayesian
The death truncation challenge is compounded by the fact that many pragmatic trials in gerontology adopt cluster randomization, either at the practice, facility, or community level, to reduce contamination, facilitate implementation, or reflect the way interventions are delivered in routine care. eldridge2012practical; murray1998design Such designs are especially common when evaluating service models, long-term care interventions, or technologies that naturally operate at a group level. For example, the Whole System Demonstrator trial of telecare randomized at the general practice level, steventon2012effect; hirani2014effect while cluster-randomized or cluster-stepped wedge designs have also been used in fall prevention, dementia care, and palliative interventions in older populations. (e.g., bhasin2020randomized; lamb2020screening Compared with individually randomized trials, CRTs introduce additional sources of correlation and heterogeneity, requiring methods that accounts for within-cluster dependence. A similar issue can arise in another type of closely related design in pragmatic trials called individually randomized group treatment trials (IRGTs), and a description of IRGT can be found elsewhere. tong2024designing Recent work has estimated SACE (and covariate conditional extensions, CSACE) in CRTs using parametric principal strata and outcome models with linear mixed effects and either Bayesian or EM-based estimation. tong2023bayesian; wang2024mixed While these approaches are interpretable and computationally convenient, they rely on linear specifications that may miss nonlinearities and high order interactions in rich baseline covariates, and they may be sensitive to functional-form misspecification.
Another common complication in pragmatic gerontologic trials is missing data. Beyond death truncation, loss to follow-up is frequent as participants relocate, withdraw consent, or become unable to complete questionnaires due to cognitive or physical decline. little2012prevention; national2011prevention In such settings, outcomes may be missing even among survivors, creating additional uncertainty in treatment effect estimation. Moreover, survival status itself can sometimes be unobserved—for example, when participants move outside the health system or registry coverage—blurring the distinction between death-related truncation and standard missingness. chatfield2005systematic These challenges are especially pronounced in aging populations, where comorbidity, frailty, and mobility constraints increase the risk of incomplete follow-up chatfield2005systematic; agogo2018longitudinal. For the estimation of treatment effect heterogeneity, failure to account for missing data can distort subgroup contrasts or mask signals of effect modification. Analyses limited to complete cases implicitly assume that attrition is independent of both treatment and prognosis, an assumption rarely plausible in gerontological research. cao2022review Consequently, principled approaches that explicitly model missingness mechanisms under reasonable data assumptions are necessary to recover credible subgroup inferences. little2012prevention; seaman2013review
This paper develops an integrated machine learning model tailored to pragmatic CRT that integrates mixed-effects BART into both the principal strata model and outcome models, while accounting for missing data in survival status and QoL outcomes. By augmenting BART with cluster-level random intercepts, the method preserves flexibility for nonlinear covariate effects and interactions while respecting within-cluster dependence dorie2022stan; spanbauer2021nonparametric; wundervald2022hierarchical. To accommodate incomplete follow-up common in pragmatic gerontologic trials, we adopt a nested missing at random scheme that treats missing survival status and, conditional on observed survival, missing outcomes as ignorable given treatment assignment and covariates tong2025bayesian. Within this framework we estimate SACE and CSACE and incorporate an exploratory, post-hoc “fit-the-fit” step that regresses individual-level CSACE estimates among likely always-survivors on baseline covariates to visualize effect modification. Our empirical application revisits the Whole Systems Demonstrator (WSD) telecare trial, focusing on EQ-5D–VAS at 12 months, and illustrates how conditional survivor causal effects can differ from the marginal intent-to-treat effect. The remainder of the paper presents the design and assumptions, the principal stratification estimands and missing data framework, the likelihood and mixed-effects BART specification, the WSD analysis and exploratory heterogeneity results, and concluding remarks.
2 Whole Systems Demonstrator Telecare Questionnaire Study
Whole Systems Demonstrator (WSD) Telecare Questionnaire Study henderson2013cost; hirani2014effect, a cluster-randomized trial conducted in England during 2008–2009. Randomization occurred at the general practice level, where a total of 1 189 participants across 204 practices in three local authorities (cluster sizes 1–26) were randomized to the telecare intervention or control with roughly a 1:1 ratio. The telecare intervention comprised in-home electronic sensors providing safety monitoring (e.g., fall and hazard alerts), while the control arm received usual care. Our primary analysis focuses on the QoL measured at 12 months by the EQ-5D–VAS index (range 0–100; higher values indicating a better quality of life). feng2014assessing Baseline characteristics including age, ethnicity, sex, highest level of education, number of comorbid conditions, living alone status, baseline deprivation score, physical health score, mental health score and EQ-5D–VAS index score by arm are summarized in Table 1. Baseline deprivation was measured with the English Index of Multiple Deprivation (IMD 2007), linked from participant postcodes to their Lower-layer Super Output Area; the IMD combines seven domains—income, employment, health and disability, education, crime, barriers to housing and services, and living environment—into an area-level score in which higher values indicate greater deprivation. noble2008english In WSD, IMD scores ranged from 1.89 to 66.49. Overall, characteristics were broadly comparable between arms: means of the continuous measures and the distributions of sex and ethnicity were closely aligned (all $p\geq 0.31$). The only notable difference was a modest shift in education levels (Cochran–Armitage $p=0.014$), with the intervention arm showing slightly higher proportions at the university/graduate categories.
Table 1: Baseline covariates in the Whole Systems Demonstrator (WSD) Telecare Questionnaire Study
Covariate Total Intervention Control $p$-value
Age, years 74.13 (13.94) 73.92 (14.32) 74.31 (13.62) 0.634
Deprivation score 28.13 (15.04) 27.70 (14.35) 28.49 (15.61) 0.363
No. of comorbid conditions 1.09 (1.45) 1.07 (1.47) 1.11 (1.44) 0.659
Physical health score 28.09 (8.62) 28.30 (8.71) 27.91 (8.55) 0.436
Mental health score 33.05 (7.87) 33.00 (7.98) 33.10 (7.78) 0.833
EQ-5D–VAS index score 53.04 (22.02) 52.77 (22.07) 53.27 (22.00) 0.697
Ethnicity 0.771
   White 1053 (88.56%) 485 (88.18%) 568 (88.89%)
   Non-white 136 (11.44%) 65 (11.82%) 71 (11.11%)
Sex 0.310
   Male 765 (64.34%) 345 (62.73%) 420 (65.73%)
   Female 424 (35.66%) 205 (37.27%) 219 (34.27%)
Highest level of education 0.014
   No formal education 779 (65.52%) 357 (64.91%) 422 (66.04%)
   GCSE/O-levels 224 (18.84%) 92 (16.73%) 132 (20.66%)
   A-levels/HNC 72 (6.06%) 30 (5.45%) 42 (6.57%)
   University level 41 (3.45%) 26 (4.73%) 15 (2.35%)
   Graduate or professional 73 (6.14%) 45 (8.18%) 28 (4.38%)
Living alone 0.603
   Yes 629 (52.90%) 286 (52.00%) 343 (53.68%)
   No 560 (47.10%) 264 (48.00%) 296 (46.32%)
  • Values are mean (SD) for continuous variables and $n$ (%) for categorical variables. $p$-values are from one-way ANOVA for continuous variables and Pearson’s $\chi^{2}$ test (Fisher’s exact test when expected cell counts <5<5) for categorical variables; for ordered education levels, the $p$-value is from the Cochran–Armitage trend test.
For the purposes of principal stratification, we classify a participant as a nonsurvivor when the 12 month status of is either death or being too ill to respond (due to dementia or loss of capacity, transition to long-term residential/nursing care, including sheltered housing, or data collection was fully reliant on a family caregiver). In the assembled data, 127 participants (10.7%) had outcomes truncated by death or serious deterioration; 62 (5.2%) were presumed alive but had unobserved outcomes for non–health-related reasons for example, some moved out of the area to non-participating general practice, and others no longer wished to remain in either the intervention or control group; and 237 (19.9%) had both survival status and the final outcome unobserved (e.g., prolonged hospitalization, refusal to share data, or equipment malfunction).
3 Methods
3.1 Notation and setup
We study a two-arm cluster-randomized trial WSD with clusters $i=1,\dots,n$, where cluster $i$ contains $N_{i}$ individuals $j=1,\dots,N_{i}$. Let $Z_{i}\in\{0,1\}$ denote the cluster-level assignment ($Z_{i}=1$ treatment; $Z_{i}=0$ control). For each individual $(i,j)$ and $z\in\{0,1\}$, $S_{ij}(z)\in\{0,1\}$ is the potential survival status at end of follow-up (11=survive, 0=death) and $Y_{ij}(z)$ is the corresponding potential non-mortality outcome, which is only well-defined when $S_{ij}(z)=1$; when $S_{ij}(z)=0$ the outcome is undefined and denoted by $\star$ zhang2003estimation. The observed data are $(Z_{i},S_{ij},Y_{ij}^{\text{obs}})$ with the usual consistency: $S_{ij}=S_{ij}(Z_{i})$ and, if $S_{ij}=1$, then $Y_{ij}^{\text{obs}}=Y_{ij}(Z_{i})$; otherwise $Y_{ij}^{\text{obs}}=\star$.
We pursue two assumptions to connect the cluster-level assignment to counterfactual outcomes. First, we invoke Stable Unit Treatment Value Assumption (SUTVA) with partial interference: for each individual $(i,j)$, the potential survival and outcome $\{S_{ij}(z),Y_{ij}(z)\}$ depend only on the assignment of their own cluster $Z_{i}$; treatment versions are unique and there is no interference across clusters, while arbitrary dependence among individuals within the same cluster is allowed. Second, we assume cluster-level randomization with positivity: for every cluster $i$, the collection $\{Y_{ij}(1),Y_{ij}(0),S_{ij}(1),S_{ij}(0)\}_{j=1}^{N_{i}}$ is independent of $Z_{i}$, and both arms occur with positive probability, $0<\Pr(Z_{i}=1)<1$. Taken together, these conditions render $Z_{i}$ ignorable for causal inference at the cluster level and underpin identification based on comparisons of observed outcomes across treatment arms.
These conditions are substantively plausible in the WSD setting. Randomization occurred at the general‐practice (cluster) level and assignment was well balanced (639 telecare vs. 550 usual care), so positivity holds and $Z_{i}$ is independent of counterfactual outcomes by design henderson2013cost; hirani2014effect. The intervention consisted of standardized, home‐installed telecare sensors delivered only to practices assigned to the telecare arm, whereas control practices continued usual care; this delivery model limits cross‐practice contamination. Any dependence among individuals within a practice (e.g., shared clinical workflows or referral patterns) is accommodated by allowing arbitrary within‐cluster correlation in our models, while interference across practices is unlikely because equipment installation and monitoring are practice‐specific and home‐based. Empirically, baseline characteristics were similar across arms (Table 1), consistent with successful randomization. Taken together, these features support SUTVA with partial interference and cluster‐level ignorability for $Z_{i}$ in the WSD trial.
3.2 Principal stratification and estimands
We use the principal stratification framework frangakis2002principal; wang2024mixed to classify each individual by the pair of potential survival indicators $\big(S_{ij}(1),S_{ij}(0)\big)$. The four strata are:
  • always-survivors $(1,1)$: survive under both treatment and control;
  • protected $(1,0)$: survive only if treated;
  • harmed $(0,1)$: survive only if assigned to control;
  • never-survivors $(0,0)$: die under both treatment and control.
For notational convenience, let $G_{ij}\in\{00,10,01,11\}$ encode principal strata membership, where $G_{ij}=ab$ corresponds to $S_{ij}(1)=a$ and $S_{ij}(0)=b$ (e.g., $G_{ij}=11$ for always-survivors, $G_{ij}=10$ for protected, $G_{ij}=01$ for harmed, $G_{ij}=00$ for never-survivors).
We adopt a monotonicity assumption tong2023bayesian; angrist1996identification stating that treatment does not reduce survival; equivalently, the harmed stratum is absent and $G_{ij}\in\{11,10,00\}$. Under this condition, observed data identify strata membership for two groups and leave two as mixtures: survivors observed under control must be always-survivors ($G_{ij}=11$), and non-survivors observed under treatment must be never-survivors ($G_{ij}=00$); by contrast, non-survivors under control comprise a mixture of protected and never-survivors ($G_{ij}\in\{10,00\}$), while survivors under treatment comprise a mixture of protected and always-survivors ($G_{ij}\in\{10,11\}$). In the WSD context, where telecare aims to improve well-being, monotonicity is substantively plausible.
Because the pair of non-mortality potential outcomes $\big(Y_{ij}(1),Y_{ij}(0)\big)$ is well-defined only for always-survivors, our primary estimand is the Survivor Average Causal Effect (SACE):
$\Delta_{\mathrm{SACE}}=\mathbb{E}\!\left[Y_{ij}(1)-Y_{ij}(0)\,\middle|\,G_{ij}=11\right].$
Let $X_{ij}$ denote baseline covariates (individual and cluster level). We also consider the Conditional SACE (CSACE),
$\Delta_{\mathrm{CSACE}}(x)=\mathbb{E}\!\left[Y_{ij}(1)-Y_{ij}(0)\,\middle|\,X_{ij}=x,\ G_{ij}=11\right],$
to characterize heterogeneity among always-survivors. In the WSD trial, we estimate both SACE and CSACE for EQ-5D-VAS.
3.3 Missing outcome and survival status
In the WSD trial, incompleteness arises from two sources: missing survival status at the end of follow-up and missing non-mortality outcomes (e.g., QoL measures). Let $R_{ij}^{S}$ and $R_{ij}^{Y}$ indicate whether survival status and the final outcome are observed, respectively ($1$ for observed and 0 for missing). When $S_{ij}=0$, the non-mortality outcome is undefined due to truncation by death, denoted $Y_{ij}=\star$. For reference, we summarize the feasible observation patterns and the recorded variables $\mathcal{O}_{ij}$ in Table 2; we will refer to each pattern by its label in bold.
Table 2: Observation patterns and recorded data for individual $(i,j)$.
Pattern (label) Recorded data $\mathcal{O}_{ij}$
Complete survivor $\{Z_{i},\ X_{ij},\ S_{ij}=1,\ R_{ij}^{S}=1,\ R_{ij}^{Y}=1,\ Y_{ij}^{\text{obs}}\}$
Death truncation $\{Z_{i},\ X_{ij},\ S_{ij}=0,\ R_{ij}^{S}=1\}$ (outcome undefined, $Y_{ij}=\star$)
Survivor, outcome missing $\{Z_{i},\ X_{ij},\ S_{ij}=1,\ R_{ij}^{S}=1,\ R_{ij}^{Y}=0\}$
Status and outcome missing $\{Z_{i},\ X_{ij},\ R_{ij}^{S}=0\}$
We formalize missingness using a nested MAR specification. Let $X_{ij}$ collect individual and cluster level covariates:
(A1) $R_{ij}^{S}\;\perp\;\big(S_{ij}(1),S_{ij}(0),Y_{ij}(1),Y_{ij}(0)\big)\;\big|\;(Z_{i},X_{ij}),$ (1)
(A2) $R_{ij}^{Y}\;\perp\;\big(Y_{ij}(1),Y_{ij}(0)\big)\;\big|\;(Z_{i},X_{ij},\ S_{ij}=1,\ R_{ij}^{S}=1).$ (2)
Assumption (1) states that, conditional on $(Z_{i},X_{ij})$, whether survival is recorded does not depend on counterfactual survivals or outcomes; Assumption (2) restricts outcome missingness to be MAR within the observed survivor subset ($S_{ij}=1$, $R_{ij}^{S}=1$). These conditions are compatible with cluster-level randomization and allow arbitrary within-cluster dependence, rendering the missingness mechanism ignorable for likelihood- or Bayesian-based inference when coupled with appropriate models for survival and outcomes.
In our analysis, for participants with observed survival but a missing final outcome we imputed the missing outcome using an arm-specific outcome model $Y_{ij}\mid(Z_{i},X_{ij},S_{ij}=1)$. When both survival status and the final outcome were unobserved we first sampling the principal stratum label $G_{ij}\in\{11,10,00\}$ from the nested Probit membership model introduced below; by monotonicity, $(G_{ij},Z_{i})$ maps to survival deterministically—survivors if $G_{ij}=11$ or if $G_{ij}=10$ and $Z_{i}=1$, and non-survivors otherwise. Those imputed as survivors then receive outcome imputation from the corresponding arm and stratum consistent model, whereas outcomes for non-survivors remain undefined ($Y_{ij}=\star$).
3.4 Likelihood and Bayesian principal stratification
We adopt a Bayesian principal stratification framework hirano2000assessing to infer principal stratum membership and to estimate SACE/CSACE. We can classify participants by observed treatment assignments and survival status at the final assessment:
  • $O(1,1)=\{(i,j)\mid Z_{i}=1,\ S_{ij}=1\}$: participants from treated clusters who survived;
  • $O(1,0)=\{(i,j)\mid Z_{i}=1,\ S_{ij}=0\}$: participants from treated clusters who did not survive;
  • $O(0,1)=\{(i,j)\mid Z_{i}=0,\ S_{ij}=1\}$: participants from control clusters who survived;
  • $O(0,0)=\{(i,j)\mid Z_{i}=0,\ S_{ij}=0\}$: participants from control clusters who did not survive.
Under monotonicity, $O(1,0)$ are identified as never-survivors ($G_{ij}=00$) and $O(0,1)$ as always-survivors ($G_{ij}=11$); by contrast, $O(1,1)$ mixes $\{11,10\}$ and $O(0,0)$ mixes $\{10,00\}$. We need to do additional modeling to due with this mixture. We use $S^{obs}$ to denote the observed survival status for all individuals and $Y^{obs}$ to denote the collection of observed non-mortality outcomes for all individuals.
Let $p_{ij,g}=\Pr(G_{ij}=g\mid X_{ij},\theta_{S})$ denote the stratum membership probabilities (summing to one across $g$). For the outcome model, define the arm and stratum specific density evaluated at the observed outcome,
$f_{ij,g,z}=f_{Y}\!\big(Y_{ij}^{\text{obs}}\mid G_{ij}=g,\ Z_{i}=z,\ X_{ij},\ \theta_{Y}\big),$
which is needed only when the outcome is well-defined in stratum $g$ under arm $z$ (i.e., $g=11$ for $z=0$; $g\in\{11,10\}$ for $z=1$). Let $\theta=(\theta_{S},\theta_{Y})$ and take independent priors $P(\theta)=P(\theta_{S})P(\theta_{Y})$.
The observed data likelihood (conditional on covariates) is
$L(Y^{\text{obs}},S^{\text{obs}},X,Z\mid\theta)$ $=\prod_{(i,j)\in O(1,1)}\Big\{\,p_{ij,11}\,f_{ij,11,1}\;+\;p_{ij,10}\,f_{ij,10,1}\,\Big\}$
$\quad\times\prod_{(i,j)\in O(1,0)}p_{ij,00}$
$\quad\times\prod_{(i,j)\in O(0,1)}p_{ij,11}\,f_{ij,11,0}$
$\quad\times\prod_{(i,j)\in O(0,0)}\big(p_{ij,10}+p_{ij,00}\big).$
This likelihood highlights where mixtures arise (exactly in $O(1,1)$ and $O(0,0)$); no outcome density is introduced for never-survivors because the non-mortality outcome is undefined. Cluster-level randomization is accommodated by conditioning on $Z_{i}$, while within-cluster dependence among individuals remains unrestricted.
3.5 Model specification
We use a nested Probit construction to impute the unobserved principal strata labels $G_{ij}\in\{00,10,11\}$. chen2024bayesian Two latent Gaussian variables are augmented for each individual: $Q_{ij}$ first separates never-survivors $(00)$ from the survivor capable strata, and $W_{ij}$ then splits the survivor capable group into protected $(10)$ versus always-survivors $(11)$.
[…]
Here $m_{Q}(\bullet)$ and $m_{W}(\bullet)$ are conditional mean functions (e.g., Probit regressions) of covariates $\mathbf{X}_{Q,ij}$ and $\mathbf{X}_{W,ij}$; they can include both individual- and cluster-level covariates. Conditional on covariates, we take $Q_{ij}$ and $W_{ij}$ as independent. This yields the strata-membership probabilities
[…]
where $\Phi(\cdot)$ is the standard normal CDF.
For outcomes, we specify arm and stratum-specific Gaussian models (defined only when the potential outcome is meaningful in that stratum/arm):
$\{\,Y_{ij}(z)\mid G_{ij}=g,\ m_{g,z}(\bullet),\ \mathbf{X}_{Y,ij}\,\}\ \sim\ \mathcal{N}\!\big(m_{g,z}(\mathbf{X}_{g,z,ij}),\ \sigma^{2}_{g,z}\big),$
with $z\in\{0,1\}$ when $g=11$ and $z=1$ when $g=10$. Concretely,
[…]
while $Y_{ij}(0)$ is undefined for $G_{ij}=10$ and both $Y_{ij}(0),Y_{ij}(1)$ are undefined for $G_{ij}=00$ by truncation-by-death. Completing the specification amounts to choosing the systematic components for the latent–Probit and outcome models, namely $m_{Q}(\mathbf{X}_{Q,ij})$, $m_{W}(\mathbf{X}_{W,ij})$, and $m_{g,z}(\mathbf{X}_{g,z,ij})$. A conventional choice is to use (generalized) linear mixed models with cluster-level random effects, which is convenient and interpretable. However, such models impose linear and additive effects (and require pre-specifying interactions), an assumption that can be too restrictive in CRT settings with potential nonlinearity and treatment–covariate interactions; misspecification may bias strata probabilities and, in turn, SACE/CSACE. Motivated by this, in what follows we adopt more flexible specifications for these mean functions—preserving random effects to account for clustering while allowing nonlinearities and high order interactions.
3.6 Integrating mixed-effects BART into principal stratification
To flexibly represent nonlinearities and higher order interactions in the conditional mean functions, we use Bayesian Additive Regression Trees (BART) chipman2010bart, which approximates a regression surface by a sum of shallow trees,
$m(\mathbf{X}_{ij})\;=\;\sum_{k=1}^{K}g_{k}\!\big(\mathbf{X}_{ij};\,\mathcal{J}_{k},\mathcal{M}_{k}\big),$ (3)
where each $g_{k}$ is a regression tree with splitting structure $\mathcal{J}_{k}$ and terminal node parameters $\mathcal{M}_{k}$. Because cluster-randomized trials exhibit within-cluster correlation, we augment each mean function with a cluster-level random intercept to absorb unobserved heterogeneity while learning covariate effects nonparametrically.
The mixed-effects BART specifications are
$m_{Q}(\mathbf{X}_{Q,ij})=\sum_{k=1}^{K}g_{Q,k}\!\big(\mathbf{X}_{Q,ij};\,\mathcal{J}_{Q,k},\mathcal{M}_{Q,k}\big)+b_{i}^{(Q)},\qquad b_{i}^{(Q)}\sim\mathcal{N}(0,\sigma_{b,Q}^{2}),$
$m_{W}(\mathbf{X}_{W,ij})=\sum_{k=1}^{K}g_{W,k}\!\big(\mathbf{X}_{W,ij};\,\mathcal{J}_{W,k},\mathcal{M}_{W,k}\big)+b_{i}^{(W)},\qquad b_{i}^{(W)}\sim\mathcal{N}(0,\sigma_{b,W}^{2}),$
$m_{g,z}(\mathbf{X}_{g,z,ij})=\sum_{k=1}^{K}g_{g,z,k}\!\big(\mathbf{X}_{g,z,ij};\mathcal{J}_{g,z,k},\mathcal{M}_{g,z,k}\big)+b_{i}^{(g,z)},\qquad(g,z)\in\{(11,0),(11,1),(10,1)\}.$
$b_{i}^{(g,z)}\sim\mathcal{N}(0,\sigma_{b,gz}^{2}).$
In practice, $m_{Q}$ and $m_{W}$ are learned by regressing the current draws of the latent membership variables on covariates using Gaussian mixed-effects BART, and $m_{11,0}$, $m_{11,1}$, $m_{10,1}$ are fit on the corresponding arm–stratum subsets; because $G_{ij}$ is latent, these subsets and the associated means are updated within the MCMC.
Mixed-effects BART has several implementations, including stan4bart dorie2022stan, mixedbart spanbauer2021nonparametric, and hebart wundervald2022hierarchical. In simulations aligned with our design, mixedbart provided a favorable balance of accuracy and computational efficiency, and we therefore use it to estimate $m_{Q}$, $m_{W}$, and $m_{g,z}$. The mixedbart model is a general framework for incorporating hierarchical random effects into BART, originally developed for longitudinal clinical trials. Their approach supports both random intercepts and slopes, though in CRTs we adopt the special case of a cluster-level intercept.
3.7 Model estimation
A fully Bayesian joint model integrating imputation within the principal stratification sampler is also feasible but beyond the scope of this illustration. Missing baseline covariates and outcomes not attributable to death/serious deterioration were completed using multiple imputation rubin1978multiple to obtain a single analysis dataset. The covariate set used in the strata membership and outcome models comprised sex, age, ethnicity, highest education level, an indicator for living alone status, number of comorbid conditions, index of multiple deprivation (IMD) score , physical and mental health scores, and the baseline EQ-5D–VAS index.hirani2014effect
We initialize the Markov chain by assigning principal–stratum labels using information identified by monotonicity: treated non-survivors are set to the never-survivor stratum and control survivors to the always-survivor stratum. For individuals in the remaining observed mixtures (treated survivors; control non-survivors), initial labels are drawn at random from the admissible strata given $(Z_{i},S_{ij})$. For units with unobserved survival, we impute their initial survival status by BART. Given these initial labels, initial estimates of outcome model $m_{g,z}(\cdot)$ can be obtained by fitting BART and initial estimate for $\sigma^{2}_{g,z}$ can be simultaneously obtained from this initial fit. For the strata–membership means $m_{Q}(\cdot)$ and $m_{W}(\cdot)$, we obtain starting values from parametric logistic regressions of the initial labels on covariates; latent normals $Q_{ij}$ and $W_{ij}$ are then drawn from the appropriate truncated Gaussian distributions implied by the initial labels. Hyperparameters for $\sigma_{g,z}^{2}$ use weakly informative inverse-gamma priors with shape and rate $a_{0}=b_{0}=0.001$ chen2024bayesian. For random intercept $\sigma_{b}^{2}$, we use the default prior in mxbart: hierarchical Inverse-Wishart with degree of freedom 3.
Each iteration updates: (i) the outcome models $m_{11,1}$, $m_{11,0}$, and $m_{10,1}$ via mixed-effects BART on the current arm–stratum subsets; (ii) the residual variances $\{\sigma^{2}_{g,z}\}$ by conjugate Gaussian updates; (iii) the strata–membership means $m_{Q}$ and $m_{W}$ via mixed-effects BART regressions of the current latent normals $Q_{ij}$ and $W_{ij}$ on covariates; (iv) the latent $Q_{ij}$ and $W_{ij}$ from their truncated normal full conditionals; (v) principal-stratum labels $G_{ij}$ using their posterior probabilities that combine the current membership model and outcome likelihood; and (vi) missing data under the nested-MAR structure: survival is sampled from the strata model, outcomes are drawn from the appropriate arm– and stratum–specific outcome model when survival is (imputed) $1$, and outcomes are undefined by design when survival is 0. We used the default settings of mixedbart (mxbart) for Gaussian responses with a random intercept per cluster.
We ran 20,000 iterations, discarding the first 10,000 as burn-in. Point estimates and 95% credible intervals for SACE and CSACE are based on posterior draws from the retained iterations. All analyses were conducted in R 4.4.1 using mixedbart and dbarts.
4 Results
Posterior inference on principal–strata composition indicated that, on average, 84.4% of participants were always-survivors, 3.2% were protected, and 12.4% were never-survivors. To summarize the CSACE, we defined the set of likely always-survivors as the union of (i) control arm participants observed to survive to 12 months and (ii) participants with a posterior probability at least 0.8 of belonging to the always-survivor stratum. Figure 1 displays each participant’s posterior probability of belonging to the always–survivor stratum plotted against the posterior mean CSACE. Most observations lie at or above 0.95 on the vertical axis, indicating high classification certainty for stratum membership. Classification uncertainty is limited and shows no systematic relation to the estimated CSACE; a small number of points have lower probabilities, but they are sparse and not concentrated at extreme CSACE values.
Figure 1: Scatter plots of the posterior probability of being in the always-survivor stratum against the posterior mean CSACE
Posterior summaries of the stratum proportions and the overall CSACE are reported in Table 3.
Table 3: Results for the Survivor Average Causal Effect Estimate and the Proportion of Recipients in Each Principal Stratum
Quantity Point estimatea 95% CrI
Proportion of always-survivors 0.844 [0.799, 0.860]
Proportion of protected 0.032 [0.000, 0.114]
Proportion of never-survivors 0.124 [0.085, 0.140]
CSACE -0.568 [3.580-3.580, $2.532$]
  • a
    Point estimate is the posterior mean. CrI: equal-tailed 95% credible interval.
Figure 2 displays the distribution of individual CSACE estimates among likely always-survivors. The left panel shows sorted posterior means with 95% CrIs; nearly all intervals cover zero, indicating limited evidence for large individual effects, although the spread of point estimates suggests heterogeneity. The right panel histogram is approximately centered at zero. The overall mean CSACE was 0.57-0.57 (95% CrI [3.58, 2.53][-3.58,\,2.53]), indicating no compelling difference between telecare and usual care in the 12 month EQ-5D–VAS summary score for the average always-survivor.
Figure 2: Distribution of CSACE on EQ-5D–VAS at 12 months among likely always-survivors. Left: sorted posterior mean CSACEs with 95% CrIs. Right: histogram of posterior mean CSACEs.
To explore effect modification among likely always-survivors, we adopted a Bayesian “fit-the-fit” strategy: hahn2020bayesian posterior mean CSACEs were used as responses in a classification and regression tree (CART) fitted to baseline covariates. The resulting tree (Figure 3) uses baseline deprivation score for the primary split at 15. Participants with deprivation $\geq 15$ had a mean CSACE of $0.61$ (95% CrI [2.66, 4.07][-2.66,\,4.07]), whereas those with deprivation <15<15 had a mean CSACE of 4.20-4.20 (95% CrI [9.23, 0.72][-9.23,\,0.72]), suggesting a greater tendency toward outcomes favoring usual care in the latter group.
Within the higher deprivation branch ($\geq 15$), a subsequent split at 50 yielded two distinct subgroups: deprivation >50>50 showed a larger mean CSACE of $4.42$ (95% CrI [2.71, 11.76][-2.71,\,11.76]), while deprivation <50<50 was near null at $0.03$ (95% CrI [3.35, 3.52][-3.35,\,3.52]). Among those most deprived (deprivation >50>50), living arrangement further refined the effect: participants not living alone exhibited the largest estimated benefit, with mean CSACE $7.13$ (95% CrI [0.88, 15.35][-0.88,\,15.35]). Within the moderate-deprivation range ($15\leq\text{deprivation}<50$), baseline mental health emerged as the key moderator: mean CSACE $1.16$ (95% CrI [3.05, 5.81][-3.05,\,5.81]) for participants with good mental health (score >32>32) versus 1.45-1.45 (95% CrI [6.22, 2.92][-6.22,\,2.92]) for poor mental health (score <32<32).
Among participants with low deprivation (<15<15), the tree highlighted the joint role of baseline mental health and baseline EQ–5D–VAS. In particular, because higher scores indicate better baseline status, those with poor baseline mental health (score <32<32) and low baseline quality of life (EQ–5D–VAS <28<28) showed a mean CSACE of 9.00-9.00 (95% CrI [19.77, 0.58][-19.77,\,0.58]), favoring usual care in that subgroup. Overall, the tree points to higher deprivation—especially in non–alone households—as associated with greater benefit from telecare, whereas low deprivation combined with poorer mental health and lower baseline EQ-5D–VAS is associated with harm. Because many terminal node intervals include zero, these subgroup findings should be interpreted as exploratory.
Figure 3: CART applied to posterior mean CSACEs for EQ-5D–VAS at 12 months among likely always-survivors. Nodes display posterior mean CSACE and 95% CrI for each subgroup; splits are on baseline covariates.
5 Discussion
This study develops and demonstrates the application of machine learning tools in identifying treatment effect heterogeneity in a pragmatic gerontological trial. Specifically, we re-examine the WSD Telecare Questionnaire Study using a principal stratification causal framework for cluster-randomized trials with outcomes truncated by death. We embed mixed-effects BART to model strata membership and arm and stratum specific outcomes, yielding estimates of the SACE and individual survival treatment effect while accounting for cluster dependence and flexible nonlinear associations between baseline covariates and outcomes. Three findings are worth highlighting.
First, the mean CSACE on EQ–5D–VAS at 12 months was close to zero, indicating no compelling average improvement among always–survivors. Because SACE conditions on principal strata rather than observed survival, this estimand differs conceptually from intent-to-treat and survivor only analyses. Our result suggests that, for individuals who would survive regardless of arm, telecare did not yield apparent improvements in overall quality of life as captured by EQ–5D–VAS.
Second, the heterogeneity analysis among likely always–survivors points to clinically meaningful variation. The most important baseline features that appear to moderate the response to telecare include community deprivation, household structure, and mental health (MS-12). Deprivation plausibly indexes environmental and social risks as well as unmet safety needs; when deprivation is higher, passive monitoring and escalation pathways have greater scope to prevent adverse events or provide reassurance, whereas in more advantaged settings the incremental value of monitoring may be limited and, for some, burdensome. Living arrangement modifies this gradient: in households with co-residents, reminders, help with equipment, interpretation of alerts, and faster response can translate monitoring into tangible benefit, while individuals living alone may not realize the same gains. Finally, poorer baseline mental health and low baseline EQ–5D–VAS mark greater symptom burden and psychosocial stress; passive monitoring alone may not address these drivers of low self-rated health and may heighten vigilance without relief, aligning with the negative signals observed in those branches. Taken together, these patterns argue for targeted deployment—prioritizing high deprivation recipients with household support while treating the subgroup signals as exploratory and requiring confirmation in pre-specified analyses. Meanwhile, it is worth highlighting that credible intervals at most terminal nodes include zero and subgroup sizes are modest; these findings should be regarded as hypothesis generating and framed as guidance for future, pre-specified confirmatory analyses rather than as definitive subgroup effects.
Third, the proposed estimation strategy handled multiple sources of incompleteness common in gerontology trials. The nested–MAR assumption separates missing survival status from, conditional on observed survival, missing outcomes, thereby permitting ignorable likelihood–based inference while retaining a clear estimand. Coupled with mixed–effects BART, the approach balances design based features of CRTs (cluster randomization, within–cluster dependence) with nonparametric flexibility in the strata and outcome models, reducing vulnerability to functional form misspecification relative to linear mixed models.
Our results relies on a series of structural assumptions. At the trial design level, we adopt SUTVA with partial interference, meaning outcomes depend only on a person’s cluster assignment, with cluster-level randomization and positivity ensuring independence of assignment and potential outcomes; in the WSD setting, where randomization occurred at the practice level and telecare was delivered within households, these conditions are reasonable, although some unobserved cross-practice interactions cannot be ruled out. At the modeling layer, we use a principal stratification framework with monotonicity for survival ($S_{ij}(1)\geq S_{ij}(0)$), which reflects the intervention’s intent, accommodate the outcomes are truncated by death, and rely on principal exchangeability, an assumption that conditional on observed covariates, models for strata membership and outcomes are sufficiently rich to capture the relevant variation. While these conditions cannot be verified directly, they are standard in mixture-based analyses and provide a defensible path to identification. For missingness, we assume a nested MAR structure, where survival status and outcomes are missing at random given treatment and covariates, which is plausible given the WSD data and reasons for follow-up loss, though we note developing methods to relax this assumption would further strengthen robustness. Finally, we recognize that the analysis of effect modification is a post hoc exploration of potential mechanism rather than a confirmatory. Nevertheless, the results highlight directions for future work and potential target population and improvement in the intervention design. Overall, our study shows that principal stratification combined with mixed effects BART yields inference that respects the trial design and is robust to nonlinear relations and higher order interactions. Substantively, the WSD reanalysis suggests that broad telecare deployment may not improve average quality of life among always–survivors; focusing implementation on participants in higher deprivation settings and those living with others may be more beneficial.
6 Conclusion
This study revisits a large, pragmatic telecare trial in older adults and asks a simple question that is often hard to answer: who is likely to benefit in terms of quality of life when some participants do not live to provide that information? By focusing on people who would be alive at follow-up and by using a cluster-aware, flexible modeling approach suited to real world trial data, we were able to look beyond an overall average and examine how responses vary across patients.
Two messages stand out. First, on average, telecare did not improve the 12 month EQ–5D–VAS among people who would survive regardless of assignment. Second, there are meaningful differences across subgroups. Signals of benefit were more apparent for participants living in more deprived settings, especially those residing with others who could support consistent engagement with the system. In contrast, participants with lower deprivation combined with poorer baseline mental health and low baseline EQ–5D–VAS showed little evidence of benefit and may do better with usual care. These patterns are exploratory, but they suggest a practical strategy: target telecare to patients and households with greater capacity to translate monitoring into action.
For clinicians and service planners, the implication is to move away from uniform deployment strategies. Prioritizing more deprived patients and ensuring household or caregiver support may yield greater gains, whereas those with substantial psychological burden and low baseline quality of life may need complementary services (e.g., mental health support, rehabilitation) alongside any monitoring technology. For researchers, future trials should prospectively evaluate such targeting rules, collect baseline factors that plausibly modify benefit, and pre-specify confirmatory subgroup analyses. The same analytical framework applies directly to other health-related quality of life outcomes collected in this study, including the SF-12 physical and mental component scores, enabling a broader view of patient-centered benefit. Replication in other settings, sensitivity analyses to key assumptions, and linkage to cost effectiveness will further clarify when and for whom telecare improves quality of life in later life.
Acknowledgments
Research in this article was partially supported by the Patient-Centered Outcomes Research Institute® (PCORI® Award ME-2020C1-19220), the United States National Institutes of Health (NIH), National Heart, Lung, and Blood Institute (grant numbers R01-HL168202, R01-HL178513), as well as by the National Institute on Aging (NIA) of the National Institutes of Health (NIH) under Award Number U54AG063546-06A1, which funds NIA Imbedded Pragmatic Alzheimer’s Disease (AD) and AD-Related Dementias Clinical Trials Collaboratory (NIA IMPACT Collaboratory), the Yale OAIC (P30AG021342), and the Yale ADRC (P30AG066508). All statements in this report, including its findings and conclusions, are solely those of the authors and do not necessarily represent the views of the NIH or PCORI® or its Board of Governors or Methodology Committee. The Whole Systems ‘Demonstrator’ (WSD) telecare trial study was funded by the Policy Research Programme in the Department of Health, UK. The views expressed are not necessarily those of the Department. The WSD study was approved by Liverpool Research Ethics Committee (ref: 08/H1005/4). We acknowledge the generous sharing of the de-identified WSD telecare trial data from Drs Shashivadan P Hirani and Stanton P Newman at City, University of London, UK.
Conflict of Interest
The authors have declared no conflict of interest.